OpenStudy (anonymous):

An architect is creating a scale drawing of a school computer lab. The length of the lab is 36 feet and the width of the lab is 48 feet. If each 12 feet of the lab equals 2 centimeters on a scale drawing, which of the following drawings is the scale drawing of the computer lab?

OpenStudy (anonymous):

no more after this i have to go

OpenStudy (anonymous):

A AND B AND C

OpenStudy (anonymous):

D

OpenStudy (anonymous):

and thats okay

OpenStudy (anonymous):

it is A bc if 12ft=2cm then 6ft=1cm so 48/6=8 then 36/6=6

OpenStudy (anonymous):

i can always help tomorrow

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

a b c or d

OpenStudy (anonymous):

i have time for one more

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

okay ill put my other one right here

OpenStudy (anonymous):

quick

OpenStudy (anonymous):

The image below is a triangle drawn inside a circle with center O: Which of the following expressions shows the area, in square inches, of the circle? (π = 3.14) 3.14 ⋅ 2 3.14 ⋅ 3 3.14 ⋅ 22 2 ⋅ 3.14 ⋅ 22

OpenStudy (anonymous):

the ones that say 22 are suppose to be 2^2

OpenStudy (anonymous):

ok its c

OpenStudy (anonymous):

A=3.14r^2

OpenStudy (mathstudent55):

\(\Large A_{circle} = \pi r^2\)

OpenStudy (study_together):

It is A because if 12m=2cm 24m=4cm 36m=6cm and the one with the number 6 in it is A so that is why it is A

OpenStudy (study_together):

You get it, right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

study together ur right but u explained wrong

OpenStudy (study_together):

What is wrong

OpenStudy (anonymous):

it is A bc if 12ft=2cm then 6ft=1cm so 48/6=8 then 36/6=6

OpenStudy (anonymous):

for the second question it is c

OpenStudy (anonymous):

@liana1026 dont forget to medal and fan me

OpenStudy (mathstudent55):

\(\Large A = \pi r^2 = 3.14 \cdot 2^2\)